[0/1 Points] DETAILS PREVIOUS ANSWERS LARCAL Find the differential dy of the given function. (Use dx for dx.) x + 2 y = 4x
Answer #1

Page No. Date : a+2 un-5 we know if U Then y = uravu 42 consider at 2 v = 42-5 = 2 Then v Put there in we get ya 1 (40-5) - 4Please like